# Quantum Monte Carlo calculations of magnetic form factors in light nuclei

### v1

@@ Abstract
The microscopic theoretical approach, describing the nucleus as a collection of nucleons interacting among each other via effective two- and three-nucleon forces, has reached remarkable sophistication [3]. Consistent comparisons with available experimental data have repeatedly confirmed the validity of this theoretical approach. Concurrent with the development of advanced computational many-body methods, the formulation of many-nucleon interactions and currents have progressed tremendously over the last three decades. Most notably, present state-of-the-art microscopic nuclear calculations adopt chiral effective field theories ($\chi$EFTs) [4] to construct many-nucleon correlations [5, 6, 7, 8, 9] and currents [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] that are grounded in the underlying theory of quantum chromodynamics (QCD) [21]. Chiral effective field theory is a low energy (and low momentum) approximation of QCD that preserves the symmetries exhibited by the fundamental theory below a breaking scale $\Lambda_{\chi}\sim 1$ GeV. Within this formalism, many-nucleon operators are arranged in a power expansion of the low-momentum scale over $\Lambda_{\chi}$. This allows for a determination of the theoretical uncertainty associated with neglecting sub-leading terms in the chiral expansion. The $\chi$EFT framework has been adopted primarily to study low-energy observables, including binding energies and spectra [22, 23, 24, 25, 26, 27, 28], electromagnetic moments [20, 29, 30, 31, 32], and transitions [33, 34, 35, 36, 37]. Analysis of the validity of $\chi$EFT at higher values of momentum transfer has been conducted principally via the study of electromagnetic form factors of the deuteron and the trinucleon systems, 3H and 3He [38, 39, 40, 41, 18, 20]. The predictability of this theory beyond the low-energy/low-momentum regime in heavier nuclei remains to be understood.
@@
We base our results on the Norfolk two- and three nucleon interactions [43, 44, 22, 45, 46] that are derived up to N3LO in the chiral expansion from a $\chi$EFT that retains pions, nucleons, and $\Delta$’s as explicit degrees of freedom. The long- and intermediate-range parts of the nucleon-nucleon interaction are modeled with one-pion and two-pion exchange mechanisms, respectively, while short-range dynamics are encoded in contact-like interactions with strengths specified by Low Energy Constants (LECs). The latter are obtained from fits to nucleon-nucleon scattering data [43, 44, 22, 45, 46, 47, 48, 49], characterized by a $\chi^{2}$ per datum close to one. Similarly, three-nucleon interactions involve pion-exchange-like mechanisms and contact terms, involving this time two LECs. In the present study, these are fit to reproduce the trinucleon biding energies, along with the triton beta decay half-life [43, 44, 22, 45, 46]. Norfolk models based on this fitting strategy are denoted in the literature with ‘NV2+3’. Here, we will discuss results based on two implementations of the ‘NV2+3’ model, namely ‘NV2+3-Ia’ and ‘NV2+3-IIb’. These models differ in the nucleon-nucleon scattering data sets adopted to constrain the two-nucleon interaction, and in the choice of regulators inherent to the theory [43, 44, 22, 45, 46].
The interaction of the external electromagnetic probe with individual nucleons and with pairs of correlated nucleons is described by one- and two-nucleon currents, respectively. The one-body current, appearing at leading order in the chiral expansion, consists of the standard convection and the spin-magnetization currents associated with an individual nucleon. In the present study, we account for additional (and negligible) relativistic corrections to this operator, which enter at N2LO in the chiral expansion. Electromagnetic two-body currents have been extensively studied within several implementations of $\chi$EFT [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Here, we adopt electromagnetic currents that are consistent with the Norfolk interactions that encompass the same nucleon-nucleon dynamics as those present in the Norfolk two-nucleon potential. These currents have been most recently summarized in Refs. [19, 20], where they are evaluated up to N3LO in the chiral expansion. Two-body electromagnetic currents consist of one- and two-pion range contributions, as well as contact terms encoding short range dynamics. They involve five unknown LECs which are determined to fits to the deuteron and trinucleon magnetic moments, as well as deuteron threshold electrodisintegration at backward angles data [20]. We note that the dominant two-nucleon current appears at NLO in the chiral expansion and consists of the well-known seagull and pion-in-flight contributions of one-pion range [50, 51, 52, 53, 54]. The NLO current is purely isovector in nature.
@@
Magnetic moments – We test the NV2+3 interaction and current models at zero value of momentum transfer by evaluating the magnetic moments of light nuclei. These calculations are summarized in Fig. 1, where the experimental data are shown by black squares, while VMC and GFMC results are shown by circles and stars, respectively. Calculations based on the one-body operator alone are indicated with empty symbols, those that include two-body currents up to N3LO are indicated with filled symbols. Pink arrows indicate the major shifts induced by the two-body currents, which are positive (negative) for neutron-rich (proton-rich) nuclei. Since the deuteron and trinucleon magnetic moments have been used to constrain the two-body current operator, predictions are solely for the $A=6-10$ nuclei. Two-body currents give a negligible correction to the LO magnetic moment in 6Li and 10B. These systems have isospin $T=0$, therefore the two-body current of one-pion range at NLO does not contribute to their magnetic moments. While the calculations at LO offer a qualitative explanation of the data, the inclusion of two-body currents brings the theory in agreement with the experiment. In 9B and 9Be, the suppression of two-body current contributions is understood by analyzing their specific structures. These systems exhibit (predominantly) a $[2\alpha+N]$ cluster structure, where $N$ is the valence nucleon, which is driving the total magnetic moment of the system. Conversely, 9C and 9Li are predominantly in the $[\alpha+^{3}{\rm He}+pp]$ and $[\alpha+^{3}{\rm H}+nn]$ cluster structures, respectively. Here, two-body currents are active within the trinucleon clusters, and between the trinucleon clusters and the valence nucleons, leading to large contributions in these systems. These findings are consistent with previous studies performed within the QMC approach [33, 29]
@@
Here, we report the first microscopic calculation of magnetic form factors of $A=6-10$ nuclei based on a $\chi$EFT formulation. Specifically, these calculations are based on the Norfolk NV2+3-IIb interaction. For this work we use the definition given in Ref. [2], i.e.,
@@
where the magnetic form factor is expressed in terms of the reduced matrix elements of the magnetic multipole operators ($M_{L}$), $J$ is the total angular momentum of the nucleus, and $q$ is the momentum transfer. We extract the multipole operators from the matrix elements of the electromagnetic current operator, as described in Ref. [42].
@@
Table 1: Summary table of experimental data sets available for the nuclei considered in this work. In the last column we emphasize if the numerical values of the magnetic form factors (or elastic cross sections) were provided in the references (N), or if we needed to digitize from the figures (D).

### v2

@@ Abstract
The microscopic theoretical approach, describing the nucleus as a collection of nucleons interacting among each other via effective two- and three-nucleon forces, has reached remarkable sophistication [3]. Consistent comparisons with available experimental data have repeatedly confirmed the validity of this theoretical approach. Concurrent with the development of advanced computational many-body methods, the formulation of many-nucleon interactions and currents have progressed tremendously over the last three decades. Most notably, present state-of-the-art microscopic nuclear calculations adopt chiral effective field theories ($\chi$EFTs) [4] to construct many-nucleon interactions [5, 6, 7, 8, 9] and currents [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] that are grounded in the underlying theory of quantum chromodynamics (QCD) [21]. Chiral effective field theory is a low energy (and low momentum) approximation of QCD that preserves the symmetries exhibited by the fundamental theory below a breaking scale $\Lambda_{\chi}\sim 1$ GeV. Within this formalism, many-nucleon operators are arranged in a power expansion of the low-momentum scale over $\Lambda_{\chi}$. This allows for a determination of the theoretical uncertainty associated with neglecting sub-leading terms in the chiral expansion. The $\chi$EFT framework has been adopted primarily to study low-energy observables, including binding energies and spectra [22, 23, 24, 25, 26, 27, 28], electromagnetic moments [20, 29, 30, 31, 32], and transitions [33, 34, 35, 36, 37]. Analysis of the validity of $\chi$EFT at higher values of momentum transfer has been conducted principally via the study of electromagnetic form factors of the deuteron and the trinucleon systems, 3H and 3He [38, 39, 40, 41, 18, 20]. The predictive power of this theory beyond the low-energy/low-momentum regime in heavier nuclei remains to be understood. Because this regime is critical to interpret experiments measuring superallowed $\beta$-decay [42, 43, 44], neutrinoless double beta decay [45, 46, 47, 48], and long-baseline neutrino oscillation experiments [49, 50, 51, 52, 53, 54, 55], it is important verify that the model is valid for these kinematics.
@@
We base our results on the Norfolk two- and three nucleon interactions [57, 58, 22, 59, 60] that are derived up to N3LO in the chiral expansion from a $\chi$EFT that retains pions, nucleons, and intermediate $\Delta$’s as explicit degrees of freedom. The long- and intermediate-range parts of the nucleon-nucleon interaction are modeled with one-pion and two-pion exchange mechanisms, respectively, while short-range dynamics are encoded in contact-like interactions with strengths specified by Low Energy Constants (LECs). The latter are obtained from fits to nucleon-nucleon scattering data [57, 58, 22, 59, 60, 61, 62, 63], characterized by a $\chi^{2}$ per datum close to one. Similarly, three-nucleon interactions involve pion-exchange-like mechanisms and contact terms, involving this time two LECs. In the present study, these are fit to reproduce the trinucleon biding energies, along with the triton beta decay half-life [57, 58, 22, 59, 60]. Norfolk models based on this fitting strategy are denoted in the literature with ‘NV2+3’. Here, we will discuss results based on two implementations of the ‘NV2+3’ model, namely ‘NV2+3-Ia’ and ‘NV2+3-IIb’. These models differ in the nucleon-nucleon scattering data sets adopted to constrain the two-nucleon interaction, and in the choice of regulators inherent to the theory [57, 58, 22, 59, 60].
The interaction of the external electromagnetic probe with individual nucleons and with pairs of correlated nucleons is described by one- and two-nucleon currents, respectively. The one-body current, appearing at leading order in the chiral expansion, consists of the standard convection and the spin-magnetization currents associated with an individual nucleon. In the present study, we account for additional (and negligible) relativistic corrections to this operator, which enter at N2LO in the chiral expansion. Electromagnetic two-body currents have been extensively studied within several implementations of $\chi$EFT [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Here, we adopt electromagnetic currents that are consistent with the Norfolk interactions that encompass the same nucleon-nucleon dynamics as those present in the Norfolk two-nucleon potential. These currents have been most recently summarized in Refs. [19, 20], where they are evaluated up to N3LO in the chiral expansion. Two-body electromagnetic currents consist of one- and two-pion range contributions, as well as contact terms encoding short range dynamics. They involve five unknown LECs which are determined to fits to the deuteron and trinucleon magnetic moments, as well as deuteron threshold electrodisintegration at backward angles [20]. We note that the dominant two-nucleon current appears at NLO in the chiral expansion and consists of the well-known seagull and pion-in-flight contributions of one-pion range [64, 65, 66, 67, 68]. The NLO current is purely isovector in nature.
@@
Magnetic moments – We test the NV2+3 interaction and current models at zero value of momentum transfer by evaluating the magnetic moments of light nuclei. These calculations are summarized in Fig. 1, where the experimental data are shown by black squares, while VMC and GFMC results are shown by circles and stars, respectively. Calculations based on the one-body operator alone are indicated with empty symbols, those that include two-body currents up to N3LO are indicated with filled symbols. Pink arrows indicate the major shifts induced by the two-body currents, which are positive (negative) for neutron-rich (proton-rich) nuclei. Since the deuteron and trinucleon magnetic moments have been used to constrain the two-body current operator, predictions are solely for the $A=6-10$ nuclei. Two-body currents give a negligible correction to the LO magnetic moment in 6Li and 10B. These systems have isospin $T=0$, therefore the two-body current of one-pion range at NLO does not contribute to their magnetic moments. While the calculations at LO offer a qualitative explanation of the data, the inclusion of two-body currents brings the theory in agreement with the experiment. In 9B and 9Be, the suppression of two-body current contributions is understood by analyzing their specific structures. These systems exhibit (predominantly) a $2\alpha+N$ cluster structure, where $N$ is the valence nucleon, which is driving the total magnetic moment of the system. Conversely, 9C and 9Li are predominantly in the $\alpha+^{3}{\rm He}+pp$ and $\alpha+^{3}{\rm H}+nn$ cluster structures, respectively. Here, two-body currents are active within the trinucleon clusters, and between the trinucleon clusters and the valence nucleons, leading to large contributions in these systems. These findings are consistent with previous studies performed within the QMC approach [33, 29]
@@
Here, we report the first microscopic calculation of magnetic form factors of $A=6-10$ nuclei based on a $\chi$EFT formulation. Because of the minimal model dependence on observables when including two-body currents through N3LO (see Ref. [73] for a detailed analysis of this dependence), we present results using the Norfolk NV2+3-IIb. For this work we use the definition given in Ref. [2], i.e.,
@@
where the magnetic form factor is expressed in terms of the reduced matrix elements of the magnetic multipole operators ($M_{L}$), $J$ is the total angular momentum of the nucleus, and $q$ is the momentum transfer. We extract the multipole operators from the matrix elements of the electromagnetic current operator, as described in Ref. [56]. Interested readers can find the explicit derivation of the relevant expressions in the companion paper of Ref. [73]. Finally, we note that the computational overhead to perform more accurate GFMC calculations provides minimal gain in interpreting the present experimental data; however, should more accurate measurements become available, it could merit a future GFMC study
@@
Table 1: Summary table of experimental data sets available for the nuclei considered in this work. In the last column we emphasize if the numerical values of the magnetic form factors (or elastic cross sections) were provided in the references (N), or if we needed to digitize from the figures (D). More details about the datasets may be found in the companion paper of Ref. [73].